# On Hardness and Approximation of Broadcasting in Sparse Graphs

### v1

@@ 4.2.1 PTAS for Double Prefix Covering
Given three sets $\mathfrak{R}_{p}(S)$, $\mathfrak{R}_{p}(C)$, and $\mathfrak{R}_{p}(D)$, there exists an algorithm that, in time $\mathcal{O}(2^{p^{3}}\,|S|\,(|S|+|C|+|D|))$, either outputs a covering of $\mathfrak{R}_{p}(S)$ using the covering sets $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$ or correctly reports that no such covering is possible.
Proof. The ground set $\mathfrak{R}_{p}(S)$ and the covering sets have constant support $p$. For any fixed distinct member of $\mathfrak{R}_{p}(S)$ like $v$ and each combination pair of members of $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$, we can encode the number of times the pair is used to cover the members of $\mathfrak{R}_{p}(S)$ with value $v$ in $\log|S|$ bits. This is because we have at most $|S|$ members with value equal to $v$.
Given that there are $p^{2}$ distinct pairs from the covering set and $p$ types of value in $\mathfrak{R}_{p}(S)$, any potential solution for the problem can be encoded in $p^{2}.p.\log|S|=p^{3}\log|S|$ bits; we can exhaustively try all and answer whether $\mathfrak{R}_{p}(C)$ can cover $\mathfrak{R}_{p}(S)$ by trying all potential solutions.
Thus, the algorithm tries $\mathcal{O}(2^{p^{3}}|S|)$ solutions. And, each solution takes linear time to check. Therefore, overall it takes $(2^{p^{3}}|S|(|S|+|C|+|D|))$ to decide whether it is possible to cover $\mathfrak{R}_{p}(S)$ using $\mathfrak{R}_{p}(C)$. $\Box$
To summarize, for a candidate covering size $m$ and ground set $\mathfrak{R}_{p}(S)$, we construct $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$ and test, via exhaustive search (Lemma 4.12), whether $\mathfrak{R}_{p}(S)$ can be covered. If no covering is found, we conclude that $\mathfrak{R}_{p}(S)$ cannot be covered with $[m]$ (bad guess). If a covering is found, we extend it to a covering with $m^{\prime}=m+\lceil m/p\rceil$, thereby obtaining a $(1+1/p)$-approximation for $\mathfrak{R}_{p}(S)$. Combining this with Proposition 4.11, we achieve an overall approximation factor of $(1+1/p)^{2}$, which approaches 1 as $p$ grows.
@@
There is a PTAS for the Double Prefix Covering problem that terminates in $\mathcal{O}(2^{p^{3}}|S|(|S|+\boldsymbol{OPT})\log\boldsymbol{OPT})$, where $p=3/\epsilon^{2}$.
Proof. Given an instance $S$ of Prefix Covering with $\epsilon>0$, set $p=\lceil 3/\epsilon^{2}\rceil$ to ensure $(1+1/p)^{2}\leq 1+\epsilon$. Construct $\mathfrak{R}_{p}(S)$ by rounding up elements in $S$. We use binary search over $m$ to find $\boldsymbol{OPT}_{p}(\mathfrak{R}_{p}(S))$, where $\boldsymbol{OPT}_{p}(S)$ is the minimum $m$ such that $S$ can be covered using $\mathfrak{R}_{p}([m])$ and $\mathfrak{R}_{p}([m-\beta])$. For each candidate set $[m]$:
@@
    Construct $\mathfrak{R}_{p}(C)$ where $C=[m]$ as per Definition 4.3.
@@
    Use exhaustive search (Lemma 4.5) to determine if $\mathfrak{R}_{p}(C)$ can cover $\mathfrak{R}_{p}(S)$ in time $\mathcal{O}(2^{p^{3}}|S|(|S|+m))$.
Therefore, the binary search routine gives the smallest $m$ such that $\mathfrak{R}_{p}(S)$ can be covered using $\mathfrak{R}_{p}([m])$ and $\mathfrak{R}_{p}([m-\beta])$. We construct a covering of $\mathfrak{R}_{p}(S)$ using $C^{\prime}=[m^{\prime}]$ and $D^{\prime}=[m^{\prime}-\beta]$, where $m^{\prime}\leq(1+1/p)m=(1+1/p)\boldsymbol{OPT}_{p}(\mathfrak{R}_{p}(S))\leq(1+1/p)\boldsymbol{OPT}(\mathfrak{R}_{p}(S))$ by 4.2. Given that $C^{\prime}=[m^{\prime}]$ and $D^{\prime}=[m^{\prime}-\beta]$ cover $\mathfrak{R}_{p}(S)$, they cover $S$ as well. Using Proposition 4.11, we have $m^{\prime}\leq(1+1/p)\boldsymbol{OPT}(\mathfrak{R}_{p}(S))\leq(1+1/p)^{2}\boldsymbol{OPT}(S)$. The running time is $\mathcal{O}(\log\boldsymbol{OPT})$ binary search iterations, each taking $\mathcal{O}(2^{p^{3}}|S|(|S|+\boldsymbol{OPT}))$ time, giving total complexity $\mathcal{O}(2^{p^{3}}|S|(|S|+\boldsymbol{OPT})\log\boldsymbol{OPT})$ where $p^{3}=\mathcal{O}(1/\epsilon^{6})$. $\Box$
Combining Lemma 4.10 and Theorem 4.13 implies the following Corollary.
@@
There exists a PTAS for Telephone Broadcasting in $k$-path graphs. In particular, for a $k$-path graph on $n$ vertices and any $\epsilon>0$, the algorithm computes a $(1+\epsilon)$-approximate broadcast schedule in time $O(2^{1/\epsilon^{6}}\,k(k+n)\log n)$.
5 Graphs of Bounded Bandwidth
@@
In this section, we present a polynomial algorithm that computes the optimal broadcasting time for graphs of bounded bandwidth. Intuitively, the bandwidth constraint limits the degree of $G$ (in contrast to other families such as $k$-cycle or $k$-path graphs), and a bandwidth ordering exposes a small “active window” of vertices at each step. This property enables a dynamic programming (DP) approach. For the remainder of this section, we assume that a bandwidth-$k$ ordering $(v_{1},\ldots,v_{n})$ of the vertices is given. For clarity of exposition, we also assume that the source of the broadcast is $v_{1}$; this restriction will be relaxed later.
@@
DP Table. We define a DP table with entries of the form $dp[i][s]$, where $i\in[0,n]$ and $s$ is a state. The value $dp[i][s]$ denotes the minimum broadcast time for informing the first $i$ vertices in the ordering while ending in configuration $s$. The state $s$ encodes all necessary information about the boundary vertices $B_{i}=\{v_{i-k+1},\ldots,v_{i}\}$, namely, how these vertices are informed and how they connect to the rest of a potential broadcast tree. Intuitively, a state describes the partial structure of a broadcast forest spanning boundary vertices together with their timing information.
@@
We fill the table row by row, starting from $i=k$ up to $i=n$. The last row corresponds to the entire graph; the minimum value among its states equals the optimal broadcast time.
Definition of states. Fix $i$, and let $u_{1},\ldots,u_{k}$ denote the boundary vertices. Let $Bf=\{x_{1},\ldots,x_{k}\}$ be the $k$ vertices immediately before $B_{i}$ and $Af=\{y_{1},\ldots,y_{k}\}$ be the $k$ vertices immediately after $B_{i}$. Together, the vertices $x_{1},\ldots,x_{k},\;u_{1},\ldots,u_{k},\;y_{1},\ldots,y_{k}$ form the active window. Every neighbor of a boundary vertex lies within this window.
@@
  • Parents. For each $u_{i}$, its parent $p(u_{i})$ (unless it is the source $v_{1}$). Each $p(u_{i})$ is chosen from at most $2k$ active vertices or a special “no parent” symbol, yielding $(2k+1)^{k}$ possibilities.
  • Ranks. For each $u_{i}$, its rank among the children of $p(u_{i})$, ordered by the rounds in which $p(u_{i})$ calls them. Since $p(u_{i})$ has at most $2k$ children, there are at most $2k$ options per vertex, and $(2k)^{k}$ total options.
  • Outgoing calls. For each $u_{i}$, the active vertex that it calls in each round after being informed. For every round slot, there are $2k$ options (one of up to 2k12k-1 children or a null option if unused), giving at most $(2k)^{2k+1}$ possibilities per $u_{i}$.
  • Inform times. The time at which the root of each subtree is informed. Since there are at most $k$ subtrees (each containing a boundary vertex), and each root may be at a round in $[n]$, this yields $n^{k}$ possibilities.
@@
If $S$ denotes the space of all states, then $|S|\;\leq\;(2k+1)^{k}\cdot(2k)^{k}\cdot(2k)^{k(2k+1)}\cdot n^{k}$, which is $\mathcal{O}(n^{k})$ for constant $k$.
@@
Filling the DP table. A state $s$ is valid for step $i$ if its parent–child assignments among nodes in its broadcast forest respect the edges of $G$. Also, the outgoing calls in the state must match the rank and parent fields. For example, if $u_{1}$ and $u_{2}$ are not adjacent in $G$ but $u_{1}$ is the parent of $u_{2}$ in $s$, then $s$ is invalid at $i$. For any invalid state at step $i$, we set $dp[i][s]=+\infty$.
Base case. When $i=k$, any valid state represents a single broadcast tree rooted at $v_{1}$ (informed at time 0). For such states, we set $dp[k][s]=\ell(s)$.
Transitions. To compute $dp[i+1][s]$ for a valid state $s$, we look for states $s^{\prime}$ that are valid at step $i$ and compatible with $s$. Compatibility means that information concerning the overlap between active states at steps $i$ and $i+1$ matches in both states: the inform times of shared vertices, their parent–child relations, and the directions and timings of edges incident to them must all agree. Further, indices shift in the natural way: for example, $u_{1}$ at step $i$ corresponds to $x_{k}$ at step $i+1$, and $u_{2}$ at step $i$ corresponds to $u_{1}$ at step $i+1$. Vertices $x_{1}$ and $y_{k}$ enter or leave the active window accordingly. See Figure 4 for an illustration.
@@
Example 9.
In Figure 4, suppose that at step $i+1$, the best compatible predecessor of state $S_{2}$ is $S_{1}$ with $dp[i][S_{1}]=120$. Since $\ell(S_{2})=109$, we obtain $dp[i+1][S_{2}]=\max\{120,109\}$.
[…]
Figure 4: An illustration of states and their compatibility for a graph of bandwidth $k=3$. The bandwidth ordering is $(\ldots,x_{1},x_{2},x_{3},u_{1},u_{2},u_{3},y_{1},y_{2},y_{3},\ldots)$. The boundary vertices $u_{i}$ are highlighted in red. When computing DP values at step $i+1$, state $S_{2}$ is compatible with state $S_{1}$. Note that the shift in the active window implies, for example, that $u_{1}$ in $S_{1}$ at step $i$ becomes $x_{3}$ in $S_{2}$ at step $i+1$, $u_{2}$ becomes $u_{1}$, $x_{2}$ becomes $x_{1}$, and so on.
Sources other than $v_{1}$. If the broadcast source is not $v_{1}$, then we consider all $O((2k)!)$ possible neighbor orderings in which the source $s$ informs its neighbors. For a fixed ordering, the problem splits into two independent subproblems: one for vertices preceding $s$ in the bandwidth ordering and one for vertices following it. The DP applies symmetrically to both sides. The only modification is in the base case: a state is valid only if it is consistent with the fixed ordering of $s$’s initial calls. Since $k$ is constant, this does not affect the asymptotic time complexity.
@@
Telephone Broadcasting can be solved optimally on graphs of bandwidth $k$ in time $f(k)\cdot n^{k^{2}+O(k)}$. In particular, for a constant $k$, there is a polynomial-time algorithm for computing an optimal broadcast schedule.
@@
Setup. Recall that we have a fixed bandwidth-$k$ ordering $(v_{1},\dots,v_{n})$. As discussed, at step $i$ the DP keeps states $s$ that summarize the broadcast restricted to the active window $W_{i}:=\{x_{1},\dots,x_{k},u_{1},\dots,u_{k},y_{1},\dots,y_{k}\}$, where $B_{i}=\{u_{1},\dots,u_{k}\}$ is the set of boundary nodes. Recall that, for a state $s$, $\ell(s)$ is the latest inform time among active vertices encoded by $s$.
Correctness. We use induction to prove the following invariant:
@@
Inductive invariant. For every $i\geq k$ and every valid state $s$ at step $i$, the entry $dp[i][s]$ of the DP table is the minimum broadcast time of any feasible broadcast scheme for informing $v_{1},\ldots,v_{i}$ which ends up at state $s$.
Base $i=k$. In the base case, the round in which each vertex of $v_{1},\ldots,v_{k}$ is informed is embedded into the state $s$. This is because they are all boundary vertices. Therefore, the minimum broadcast time over schedules realizing $s$ is exactly $\ell(s)$, so $dp[k][s]=\ell(s)$ as set by the algorithm.
Step $i\to i+1$. Let $s$ be any valid state at step $i+1$. Choose a compatible predecessor $s^{*}$ at step $i$ that minimizes $dp[i][s^{*}]$.
@@
By the inductive hypothesis there is broadcast scheme $\mathcal{S}^{*}$ that realizes $s^{*}$ and informs $v_{1},\ldots,v_{i}$ in $dp[i][s^{*}]$ rounds. Extend $\mathcal{S}^{*}$ inside the active window to realize $s$ at step i+1i{+}1 (this is possible by compatibility and validity). The broadcast time of the extended scheme is $\max(dp[i][s^{*}],\ell(s))$, which is exactly the value assigned by the recurrence. Hence, dp[i+1][s]dp[i{+}1][s] upper-bounds the true optimum among schedules realizing $s$.

### v2

@@ 4.2 EPTAS for Melon Graphs
Given three sets $\mathfrak{R}_{p}(S)$, $\mathfrak{R}_{p}(C)$, and $\mathfrak{R}_{p}(D)$, there exists an algorithm that, in time $p^{\mathcal{O}(p^{3})}\cdot\operatorname{poly}(|S|+|C|+|D|)$, either outputs a covering of $\mathfrak{R}_{p}(S)$ using the covering sets $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$ or correctly reports that no such covering is possible.
Proof. Let $V$, $T_{C}$, and $T_{D}$ be the set of distinct elements in $\mathfrak{R}_{p}(S)$, $\mathfrak{R}_{p}(C)$, and $\mathfrak{R}_{p}(D)$, respectively. Then $|V|\leq p$ and $|T_{C}|,|T_{D}|\leq p$. For each tuple $(i,j)\in(T_{C}\cup\{\emptyset\})\times(T_{D}\cup\{\emptyset\})$ (except $(\emptyset,\emptyset)$) and each ground type $v\in V$, introduce an integer variable $x_{i,j,v}\in\mathbb{Z}_{\geq 0}$, intended to denote how many times the pair $(i,j)$ is used to cover elements of type $v$. There are $((|T_{C}|+1)(|T_{D}|+1)-1)|V|\leq p^{3}$ variables.
We write a feasibility ILP whose constraints enforce that (i) every ground element is covered exactly once, and (ii) cover-items are not used more than their multiplicities, (iii) every ground element is covered with the right set of values.
Cover all ground elements. For each $v\in V$, we let $\mu_{S}(v)$ be the multiplicity of $v$, that is, $\mu_{S}(v):=|\{s\in\mathfrak{R}_{p}(S):s=v\}|$. To ensure $v$ is covered exactly once, we add a constraint:
$\sum_{(i,j)\in(T_{C}\cup\{\emptyset\})\times(T_{D}\cup\{\emptyset\})}x_{i,j,v}\;=\;\mu_{S}(v).$ (3)
Do not overuse cover-items. For each $t\in T_{C}$, let $\mu_{C}(t)$ be the multiplicity of type $t$ in $\mathfrak{R}_{p}(C)$, that is, $\mu_{C}(t):=|\{s\in\mathfrak{R}_{p}(C):s=t\}|$. Similarly, define $\mu_{D}(t)$ for each $t\in T_{D}$. Each time we use a pair $(i,j)$ we consume one item of type $i$ and one item of type $j$ (with the understanding that if $i=j$ then we consume two copies). Thus, for each $t\in T_{C}$ we impose
$\sum_{v\in V}\ \sum_{j\in T_{D}\cup\{\emptyset\}}\,x_{t,j,v}\;\leq\;\mu_{C}(t).$ (4)
Similarly, we impose the same constraint for $\mu_{D}(t)$ for each $t\in T_{D}$:
$\sum_{v\in V}\ \sum_{i\in T_{C}\cup\{\emptyset\}}\,x_{i,t,v}\;\leq\;\mu_{D}(t).$ (5)
Type-feasibility. Finally, we forbid pairs that cannot cover type $v$ by setting $x_{i,j,v}=0$ whenever $(i,j)$ is not a valid set for covering $v$, that is, when the sum of $i$ and $j$ (if it exists) is less than $v$.
This ILP has at most $p^{3}$ integer variables, and the number of constraints is $\mathcal{O}(p^{3})$.
By Lenstra’s theorem [22] (and subsequent improvements by Kannan [21]), integer linear feasibility with a fixed number $d$ of variables can be solved in time $d^{\mathcal{O}(d)}\cdot\operatorname{poly}(L)$, where $L$ is the encoding length of the ILP. Here $d\leq p^{3}$ and $L=\operatorname{poly}(|S|+|C|+|D|+p)$, hence the feasibility test runs in $p^{\mathcal{O}(p^{3})}\cdot\operatorname{poly}(|S|+|C|+|D|)$ time. If feasible, the resulting integer solution specifies a valid covering; otherwise, none exists. $\Box$
To summarize, for a candidate covering size $m$ and ground set $\mathfrak{R}_{p}(S)$, we construct $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$ and test, via a feasibility ILP (Lemma 4.12), whether $\mathfrak{R}_{p}(S)$ can be covered. If no covering is found, we conclude that $\mathfrak{R}_{p}(S)$ cannot be covered with $[m]$ (bad guess). If a covering is found, we extend it to a covering with $m^{\prime}=m+\lceil m/p\rceil$, thereby obtaining a $(1+1/p)$-approximation for $\mathfrak{R}_{p}(S)$. Combining this with Proposition 4.11, we achieve an overall approximation factor of $(1+1/p)^{2}$, which approaches 1 as $p$ grows.
@@
There is a $(1+\epsilon)$-approximation for the Double Prefix Covering problem that terminates in $\mathcal{O}(p^{\mathcal{O}(p^{3})}poly(|S|,\boldsymbol{OPT}))$, where $p=3/\epsilon^{2}$.
Proof. Given an instance $S$ of Double Prefix Covering with $\epsilon>0$, set $p=\lceil 3/\epsilon^{2}\rceil$ to ensure $(1+1/p)^{2}\leq 1+\epsilon$. Construct $\mathfrak{R}_{p}(S)$ by rounding up elements in $S$. We use binary search over $m$ to find $\boldsymbol{OPT}_{p}(\mathfrak{R}_{p}(S))$, where $\boldsymbol{OPT}_{p}(S)$ is the minimum $m$ such that $S$ can be covered using $\mathfrak{R}_{p}([m])$ and $\mathfrak{R}_{p}([m-\beta])$. For each candidate set $[m]$:
@@
    Construct $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$ where $C=[m],D=[m-\beta]$ as per Definition 4.3.
@@
    Use the feasibility ILP (Lemma 4.12) to determine if $\mathfrak{R}_{p}(C)$ and $\mathfrak{R}_{p}(D)$ can cover $\mathfrak{R}_{p}(S)$ in time $\mathcal{O}(p^{O(p^{3})}poly(|S|,{\boldsymbol{br}^{*}})$.
Therefore, the binary search routine gives the smallest $m$ such that $\mathfrak{R}_{p}(S)$ can be covered using $\mathfrak{R}_{p}([m])$ and $\mathfrak{R}_{p}([m-\beta])$. We construct a covering of $\mathfrak{R}_{p}(S)$ using $C^{\prime}=[m^{\prime}]$ and $D^{\prime}=[m^{\prime}-\beta]$, where $m^{\prime}\leq(1+1/p)m=(1+1/p)\boldsymbol{OPT}_{p}(\mathfrak{R}_{p}(S))\leq(1+1/p)\boldsymbol{OPT}(\mathfrak{R}_{p}(S))$ by 4.2. Given that $C^{\prime}=[m^{\prime}]$ and $D^{\prime}=[m^{\prime}-\beta]$ cover $\mathfrak{R}_{p}(S)$, they cover $S$ as well. Using Proposition 4.11, we have $m^{\prime}\leq(1+1/p)\boldsymbol{OPT}(\mathfrak{R}_{p}(S))\leq(1+1/p)^{2}\boldsymbol{OPT}(S)$. The running time is $\mathcal{O}(\log\boldsymbol{OPT})$ binary search iterations, each taking $\mathcal{O}(p^{\mathcal{O}(p^{3})}poly(|S|,{\boldsymbol{br}^{*}}))$ time, giving total complexity $p^{\mathcal{O}(p^{3})}poly(|S|,{\boldsymbol{br}^{*}}))$ where $p^{3}=\mathcal{O}(1/\epsilon^{6})$. $\Box$
EPTAS for Telephone Broadcasting in Melon graphs
Given the equivalence between Telephone Broadcasting in melon graphs and the Double Prefix Covering problem, as established in Lemma 4.10 along with the EPTAS presented in Theorem 4.13 for the Double Prefix Covering problem, we can conclude the following.
@@
There exists an EPTAS for Telephone Broadcasting in melon graphs. In particular, for a melon graph on $n$ vertices and any $\epsilon>0$, the algorithm computes a $(1+\epsilon)$-approximate broadcast schedule in time $O(2^{1/\epsilon^{6}}\,k(k+n)\log n)$.
Multi-source Broadcasting in Melon Graphs
The EPTAS of Corollary 4.14 extends naturally to the multi-source broadcasting setting. Suppose we are given an instance $I=(G,S)$ of multi-source broadcasting in $G$, where $S$ is the set of all sources, along with a fixed parameter $\epsilon>0$. Let $\boldsymbol{br}$ be a broadcast-time guess value. We describe a procedure that, when $\boldsymbol{br}={\boldsymbol{br}^{*}}(G,S)$, outputs a valid broadcasting scheme that completes within $(1+\epsilon)\boldsymbol{br}+1$. Otherwise, it returns either Bad-Guess or a scheme with no guarantee on its broadcast time. As we do not know the optimal broadcast time, we iterate over all $\boldsymbol{br}\in[\lceil\log n\rceil,n]$ and take the scheme with the best broadcasting time among all returned schemes. This way, we can achieve broadcast time $(1+\epsilon){\boldsymbol{br}^{*}}(I)+1$.
Preprocessing. For each non-endpoint source $s_{i}$ in the melon graph, we remove all vertices that lie within distance $\boldsymbol{br}$ of $s_{i}$ on their respective paths, excluding the $s$ and $t$. (The vertex $s_{i}$ is also removed.) If the endpoints are not initially source vertices, let $s_{ms}$ and $s_{mt}$ be the sources closest to $s$ and $t$, respectively, at distance $d_{\min}(s)$ and $d_{\min}(t)$.
Let $G^{\prime}$ denote the remaining graph and $d_{\min}=\min(d_{\min}(s),d_{\min}(t))$. If $G^{\prime}$ is disconnected, or if $\max(d_{\min}(s),d_{\min}(t))>\boldsymbol{br}$, then there exists a vertex whose distance from every source exceeds $\boldsymbol{br}$, and we correctly return Bad-Guess.
The preprocessing removes a window of at most $\boldsymbol{br}$ vertices around each non-center source. Note that all removed vertices can be informed within $\boldsymbol{br}+1$ rounds by broadcasting from their respective sources.
Reduction to covering. Consider an instance $I^{\prime}=(G^{\prime},\{s,t\},\alpha)$ of Telephone Broadcasting, where $s$ receives the message at time 0 and $t$ at time $\alpha$. We associate $I^{\prime}$ with an instance $\mathcal{S}$ of the Double Prefix Covering problem, defined analogously to Lemma 4.10 for melon graphs. The only difference is that some elements of the ground multiset $\mathcal{S}$, which we refer to as partial path elements, correspond to partially-covered paths in $G^{\prime}$ rather than full paths. Each such path, depending on the endpoint that it is connected to, must be covered by exactly one of the endpoints (it cannot be covered by both covering sets).
This modified covering instance can be solved in the same way as discussed in Section 4.2. In particular, the rounding steps remain unchanged. The only additional constraint arises in the final step (Lemma 4.12), where we ensure that each partial path element in the rounded ground set is covered by its corresponding rounded cover set. This produces a broadcast schedule for $I^{\prime}$ that completes within $(1+\epsilon)\,{\boldsymbol{br}^{*}}(I^{\prime})$ rounds.
Combining the solutions. Overall, broadcasting in the original multi-source instance $G$ completes within
$\max\!\left\{\,\boldsymbol{br}+1,\ d_{\min}+1+(1+\epsilon)\,{\boldsymbol{br}^{*}}(I^{\prime})\,\right\}.$
The term $\boldsymbol{br}+1$ accounts for informing all vertices removed during the preprocessing step. The second term captures the time required to inform the closest endpoint and possibly $1$ extra round if the same source covers both endpoints (one of the endpoints can be informed through the other endpoint), which is $d_{\min}+1$ rounds, followed by the execution of the single-source broadcast within $I^{\prime}$.
Suppose $br={\boldsymbol{br}^{*}}(I)$. Now, if $G^{\prime}$ has no vertices other than $s$ and $t$, then we have found a broadcasting schedule with $\boldsymbol{br}+1={\boldsymbol{br}^{*}}(I)+1$ rounds. Otherwise, in the optimal scheme, vertices in $G^{\prime}$ must be informed through the $s$ and $t$. This is because the vertices in $G^{\prime}$ are the ones that were not removed when forming $G^{\prime}$ and thus are at a distance further than $\boldsymbol{br}={\boldsymbol{br}^{*}}(I)$ from the sources in their respective cycles. Therefore, as the center cannot be informed earlier than round $d_{\min}$, we can write ${\boldsymbol{br}^{*}}(I)\geq d_{\min}+{\boldsymbol{br}^{*}}(I^{\prime})$. Hence, the broadcast time of our scheme is at most
$\max\!\left\{\,\boldsymbol{br}+1,\ d_{\min}+1+(1+\epsilon)\,{\boldsymbol{br}^{*}}(G^{\prime},s)\,\right\}\leq(1+\epsilon){\boldsymbol{br}^{*}}(G,S)+1.$
5 Graphs of Bounded Cutwidth
@@
In this section, we give a polynomial-time algorithm for optimally solving the multi-source broadcasting problem on graphs of bounded cutwidth. Intuitively, the cutwidth constraint limits the degree of $G$ (in contrast to other families such as cycle-star or melon graphs), and a cutwidth ordering exposes a small “active window” of vertices at each step. This property enables a dynamic programming (DP) approach. We assume throughout that a cutwidth-$k$ ordering $(v_{1},\ldots,v_{n})$ is given; all references to “before” and “after” are with respect to this ordering. Recall that such an ordering can be computed in fixed-parameter tractable time for constant $k$ [28].
@@
DP Table. We define a DP table with entries of the form $dp[i][s]$, where $i\in[0,n]$ and $s$ is a state. The value $dp[i][s]$ denotes the minimum broadcast time for informing the first $i$ vertices in the cutwidth ordering while ending in state $s$. Define the set of boundary vertices as $B_{i}=\{v_{j}|j\leq i,v_{j}\text{ has at least a neighbor after }v_{i}\}$. Given that cutwidth is $k$, it holds that $|B_{i}|\leq k$. The state $s$ encodes all necessary information about the boundary vertices, namely, how these vertices are informed and how the vertices appearing after $v_{i}$ can connect to them to form a potential broadcast tree.
@@
Intuitively, a state describes the partial structure of a broadcast forest spanning boundary vertices together with their timing information. We fill out the table row by row, starting with $i=0$ and ending with $i=n$. The last row corresponds to the entire graph, and the minimum value among its states indicates the optimal broadcast time.
Definition of states. Fix $i$, and recall $B_{i}$ denotes the set of boundary vertices. We define the active window of $i$, denoted by $A_{i}$, as a set formed by the union of $B_{i}$ and all neighbors of vertices in $B_{i}$. Given a cutwidth-$k$ ordering, the degree of every vertex is at most $2k$. For an index $i$, the active window contains all vertices in $B_{i}$ (at most $k$ vertices) and their neighbors. Therefore, its total size is bounded by $k+2k^{2}$.
@@
  • Parents. For each $u\in B_{i}$, its parent $p(u)$ (unless it is a source vertex). Each $p(u)$ is chosen from at most $2k$ active vertices or a special “no parent” symbol for the sources, yielding $(2k+1)^{k}$ possibilities.
  • Ranks. For each $u\in B_{i}$, its rank among the children of $p(u)$, ordered by the rounds in which $p(u)$ calls them. Since $p(u)$ has at most $2k$ children, there are at most $2k$ options per vertex, and $(2k)^{k}$ total options.
  • Outgoing calls. For each $u\in B_{i}$, the active vertex that it calls in each round after being informed. For every round slot, there are $2k+1$ options (one of up to $2k$ children or a null option if the slot is unused), giving at most $(2k+1)^{2k+1}$ possibilities per $u\in B_{i}$.
  • Inform times. The time at which the root of each subtree is informed. Since there are at most $k$ subtrees (each containing a boundary vertex), and each root may be informed at a round in $[n]$, this yields $n^{k}$ possibilities.
@@
If $S$ denotes the space of all states, then
$|S|$ $\;\leq\;f(k)\cdot n^{k},$
$\text{\ where \ \ }f(k)$ $=(2k+1)^{k}\cdot(2k)^{k}\cdot(2k+1)^{k(2k+1)}.$
In particular, $|S|=\mathcal{O}(n^{k})$ for constant $k$.
@@
Filling Out the DP table. A state $s$ is valid for step $i$ if its parent–child assignments among nodes in its broadcast forest respect the edges of $G$. For example, if $u_{1}$ and $u_{2}$ are not adjacent in $G$ but $u_{1}$ is the parent of $u_{2}$ in $s$, then $s$ is invalid at $i$. Furthermore, the inform times of all source vertices must be 0. For any invalid state at step $i$, we set $dp[i][s]=+\infty$.
Base case. When $i=0$, the active window is an empty set. Therefore, there is only one null valid state $s_{\phi}$, for which $dp[0][s_{\phi}]=0$.
Transitions. To compute $dp[i+1][s]$ for a valid state $s$, we check states like $s^{\prime}$ that are valid at step $i$ and compatible with $s$. Compatibility means that information concerning the overlap between active states at steps $i$ and $i+1$ matches in both states: the inform times of shared vertices, their parent–child relations, and the directions and timings of edges incident to them must all agree.
@@
Telephone Broadcasting can be solved optimally on graphs of cutwidth $k$ in time $f(k)^{2}\cdot n^{2k+1}$. In particular, for a constant $k$, there is a polynomial-time algorithm for computing an optimal broadcast schedule.
@@
Setup. Recall that we have a fixed cutwidth-$k$ ordering $(v_{1},\dots,v_{n})$. As discussed, at step $i$, the DP table has states $s$ that summarize the broadcast restricted to the active window $A_{i}$ with boundary set $B_{i}$. Recall that, for a state $s$, $\ell(s)$ is the latest inform time among active vertices encoded by $s$.
Correctness. Define the overlap signature of a state $s$ at step $i$ as the restriction $\pi_{i}(s)$ of $s$ to the overlap $A_{i}\cap A_{i+1}$: this includes (i) parent/child relations among vertices in the overlap, (ii) the directions/times of incident used edges, and (iii) their inform times. By definition, two states $s^{\prime}$ at $i$ and $s$ at i+1i{+}1 are compatible iff $\pi_{i}(s^{\prime})=\pi_{i}(s)$. We use induction to prove the following invariant:
@@
Claim 1 (Inductive invariant). For every $i\geq 0$ and every valid state $s$ at step $i$, the entry $dp[i][s]$ of the DP table is the minimum broadcast time of any feasible broadcast scheme for informing $v_{1},\ldots,v_{i}$ which ends up at state $s$ for the active set.
Base $i=0$. As there are no vertices in this case, the answer for state $s_{\phi}$ is trivially 0 to inform an empty set.
@@
Step $i\to i+1$. Let $s$ be any valid state at step $i+1$. Choose a compatible predecessor $s^{*}$ at step $i$ that minimizes $dp[i][s^{*}]$. By the inductive hypothesis there is broadcast scheme $\mathcal{S}^{*}$ that realizes $s^{*}$ and informs $v_{1},\ldots,v_{i}$ in $dp[i][s^{*}]$ rounds. Extend $\mathcal{S}^{*}$ inside the active window to realize $s$ at step i+1i{+}1 (this is possible by compatibility of $s$ and $s^{*}$ and their validity). The broadcast time of the extended scheme is $\max\{dp[i][s^{*}],\ell(s)\}$, which is exactly the value assigned by the recurrence. Hence, dp[i+1][s]dp[i{+}1][s] upper-bounds the true optimum among schedules realizing $s$.